compounding periods example:
the investment is paid at 8%, compounded quarterly, for 2 years, 1,000 investment and its future value 1171.66.
so i took {1+ 0.08}^4= (1.04)^4-1171.66 (((((cant type the problem right
over 4 on this screen)))))
anyway, can these formulas be done on a traditional 10 key adding machine or do i need a different calculator. i can not arrive at this answer.
i have multiplied, divided, subtracted, hit the wall, cursed, slapped the fridge, i just can not get the answer to come out and i think i know how to but i think i am missing a step on the term and power calc. i could sure use an explanation on how to properly arrive at the answer. thanks.

Found 2 solutions by scott8148, rfer:
Answer by scott8148(6628)

you can get it without the power function ___ you just have to multiply several times

8% compounded quarterly for two years is 8 compoundings (2 * 4) at 2% each (8% / 4)

(1 + .02)^8 = 1.171659...

Answer by rfer(16322)

It really helps if you have a Scientific or Business Calculator that you can perform funtions with exponents. Get a Texas Instrument for about $15.
FV=P(1+r/n)^nt
FV=1000(1+.08/4)^4*2
FV=1000(1+.02)^8
FV=1000(1.02)^8
FV=1000*1.171659
FV=1171.66